Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

If a person purchases 15 of the 3,000 tickets sold in a raffle that awards one prize, what is the probability that...

Expert replies
by BTGmoderatorLU » Thu Jan 14, 2021 10:07 am
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

Source: GMAT Prep

If a person purchases 15 of the 3,000 tickets sold in a raffle that awards one prize, what is the probability that this person will not win?

A. 0
B. 1/200
C. 1/2
D. 199/200
E. 1

The OA is D
Join the discussion
Source: — Problem Solving |

BTGmoderatorLU wrote: ↑
Thu Jan 14, 2021 10:07 am
Source: GMAT Prep

If a person purchases 15 of the 3,000 tickets sold in a raffle that awards one prize, what is the probability that this person will not win?

A. 0
B. 1/200
C. 1/2
D. 199/200
E. 1

The OA is D
Solution:

The probability of not winning is (3,000 - 15) / 3,000 = 2,985 / 3,000 = 597 / 600 = 199 / 200.

Alternate solution:

The probability of winning is 15/3,000 = 1/200; thus, the probability of not winning is 1 - 1/200 = 199/200.

Answer: D

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion

BTGmoderatorLU wrote: ↑
Thu Jan 14, 2021 10:07 am
Source: GMAT Prep

If a person purchases 15 of the 3,000 tickets sold in a raffle that awards one prize, what is the probability that this person will not win?

A. 0
B. 1/200
C. 1/2
D. 199/200
E. 1

The OA is D
Let's use the complement: P(Event A happening) = 1 - P(Event A not happening)

Given: P(person wins) [m]= [fraction]15/3000[/fraction][/m]

So, P(person does NOT win) [m]= 1 - [fraction]15/3000[/fraction][/m]

Simplify to get: P(person does NOT win) [m]= 1 - [fraction]1/200[/fraction][/m]

Rewrite with common denominator to get: P(person does NOT win) [m]= [fraction]200/200[/fraction] - [fraction]1/200[/fraction] = [fraction]199/200[/fraction][/m]

Answer: D

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion